OpenStudy (anonymous):

medal-----I just need help figuring out this question not just the answer please The table below represents the velocity of a car as a function of time: Time Velocity (hour) (miles/hours) x y 0 50 1 52 2 54 3 56 Part A: What is the y-intercept of the function, and what does this tell you about the car? Part B: Calculate the average rate of change of the function represented by the table between x = 1 to x = 3 hours, and tell what the average rate represents. Part C: What would be the domain of the function if the velocity of the car was measured until it reached 60

OpenStudy (anonymous):

Help please

OpenStudy (anonymous):

Could you please help me if you can @Nnesha or @whpalmer4 or if you know someone you can tag to help

OpenStudy (whpalmer4):

What is the definition of the \(y\)-intercept?

OpenStudy (anonymous):

yes here

OpenStudy (whpalmer4):

So, what is the definition of the \(y\)-intercept?

OpenStudy (anonymous):

it is the point on a graph when the x is equalled to 0....?

OpenStudy (whpalmer4):

Well, it's the value of \(y\) at the point on the graph where \(x=0\). Also the point where the graph crosses the \(y\)-axis.

OpenStudy (whpalmer4):

So if we graphed the data in the table, would the \(x\)-axis be time, or velocity?

OpenStudy (anonymous):

x wold be time because the velocity would be y

OpenStudy (whpalmer4):

good. so now you can tell me the answer to part A...

OpenStudy (anonymous):

so when x is 0 y is 50

OpenStudy (whpalmer4):

Yes. And as \(x\) represents time, that's the initial velocity of the car, right? 50 miles/hour

OpenStudy (anonymous):

yes so all i have to write would be.... the y intercept would be 50 if x is 0 im not sure how to word it

OpenStudy (whpalmer4):

How about "the y-intercept of the function is 50 miles/hour"?

OpenStudy (anonymous):

ok ..... :) but it also says what does this tell you about the car? i dont know how to figure that out or does that answer the entire question ? @whpalmer4

OpenStudy (whpalmer4):

If you look back a few responses, I think you will find that I made a comment about what that represents...

OpenStudy (whpalmer4):

The data in the table shows the car starting at 50 mph and going faster with each passing hour, right?

OpenStudy (anonymous):

yes by 2 miles per hour

OpenStudy (whpalmer4):

no, but close: it changes by 2 miles per hour per hour...

OpenStudy (whpalmer4):

or miles/hour^2

OpenStudy (anonymous):

yes :) your right

OpenStudy (whpalmer4):

no, "you're right" :-)

OpenStudy (whpalmer4):

On to part B! How do you find the average rate of change?

OpenStudy (anonymous):

average rate of change for an function f(x) is intervals from a to b so the formula is f(b)-f(a) over b-a

OpenStudy (whpalmer4):

okay, do you understand how to translate your table values into a,b,f()?

OpenStudy (anonymous):

@whpalmer4 i was reading my notes from class would F(a) be x and f (b) be y

OpenStudy (anonymous):

so would it be from the table 52-50 over 1-0 ??? not sure

OpenStudy (whpalmer4):

Well, \(x\) takes on the values of \(a\) and \(b\) and \(y = f(x)\) So at \(x=1 \text{ hour}\) we have \(a = x = 1\text{ hour}\) and \(f(1) = \) what?

OpenStudy (anonymous):

52

OpenStudy (whpalmer4):

Yes. so \(a = 1\) and \(f(a) = 52\) Now how about \(b\) and \(f(b)\)?

OpenStudy (anonymous):

b =2 and f(b)=2

OpenStudy (anonymous):

f(b) =54 oops

OpenStudy (whpalmer4):

why does b = 2? and f(b) is the value of the function at x=b=2, which doesn't appear to be 2 when I read the table... Here's the initial problem again: Part B: Calculate the average rate of change of the function represented by the table between x = 1 to x = 3 hours, and tell what the average rate represents. ^^^^^^^^^

OpenStudy (whpalmer4):

well, my attempt to underline "\(x=3\)" didn't work so well!

OpenStudy (anonymous):

ohhh So b =3 and f(b)=56

OpenStudy (whpalmer4):

yes! we are trying to find the average rate of change over that 2 hour span, so \(a\) is the value of \(x\) or \(t\) at the start, and \(b\) is the value of \(x\) or \(t\) at the end, and \(f(a)\) and \(f(b)\) are the values from the table.

OpenStudy (whpalmer4):

So our average rate of change between \(a\) and \(b\) is as your formula states: \[\text{avg rate of change} = \frac{f(b)-f(a)}{b-a}\]

OpenStudy (whpalmer4):

Can you plug in the values?

OpenStudy (anonymous):

56-52/ 3-1=4/2 =2

OpenStudy (whpalmer4):

Okay, if you are going to write equations on one line like that, you MUST use parentheses appropriately to show what isn't being shown by position! what you wrote is actually equal to this: \[56 - \frac{52}{3} - 1 = \frac{4}{2}\] I know that isn't what you meant. (56-52)/(3-1) = 4/2 would unambiguously convey the intended meaning

OpenStudy (whpalmer4):

What will the units for your answer be? miles? miles per hour? something else?

OpenStudy (anonymous):

oh ok thanks for that tip and my units would be miles per hour or miles/hour

OpenStudy (anonymous):

so ithe change of rate would be 2 miles per hour

OpenStudy (whpalmer4):

right number, wrong unit \[\frac{56 \text{ miles/hr} - 52 \text{ miles/hr}}{3\text{ hr} - 1\text{ hr}} = \frac{4\text{ miles/hr}}{2\text{ hr}} = 2 \text{miles/hr}^2\] \]

OpenStudy (anonymous):

can you explain why is it hr^2

OpenStudy (anonymous):

I dont know how to get the little 2 sorry

OpenStudy (whpalmer4):

right? if we continue accelerating at this rate, for every hour that passes, we will increase our speed by 2 miles/hour. If we accelerate for 10 hours, we will add \[10\text{ hours}*2\frac{\text{ miles}}{\text{hour*hour}} = 10\cancel{\text{ hours}}*2\frac{\text{ miles}}{\text{hour}*\cancel{\text{hour}}} \] \[= 20 \frac{\text{ miles}}{\text{hour}}\] to our speed

OpenStudy (anonymous):

ohhh yes ........ :0

OpenStudy (anonymous):

thanks for explaining it

OpenStudy (anonymous):

my mom says we have to go to church can i see if your on later to finish :)

OpenStudy (whpalmer4):

the way it "builds up" in physics terminology: \[x = v t\]distance = speed * time \[v = a t\]velocity = acceleration * time

OpenStudy (whpalmer4):

Sure, say hi to the big guy for me :-)

OpenStudy (anonymous):

lol i will thank you so much